Calculate Return on Assets Ratio:

2018 2017
Net Income 15297 1300
Total Assets 152954 157303
Answer #1

Return on Assets Ratio:

= Net Income ÷ Average Total Assets

1) For 2018:

=15297 ÷( 152954 + 157303/2)

= 15297 ÷ 155128.5

= 9.86%

2) For 2017 avg total asset is not availble.

Note:

If averge asset is not available than we can also calculate ratio with the help of total asset.

Return on Asset:

Net income / Total Assets = ROA
2018 15297 / 152954 = 10%
2017 1300 / 157303 = 0.83%